# A Visual Perception-Based Tunable Framework and Evaluation Benchmark for H.265/HEVC ROI Encryption

### v1

@@ V ROI Encryption Evaluation Benchmark
To address this issue, We pioneer a new comprehensive ROI encryption evaluation benchmark to provide a standardized evaluation platform for different ROI encryption algorithms. We believe that the performance of a ROI encryption algorithm depends on two key factors: First, the IoU between the encrypted regions and the detected ROIs is used to quantify how accurately sensitive areas are covered by the encryption. Secondly, the perturbation effect within the extracted ROI regions, which affects whether the content in these regions is sufficiently perturbed to become unrecognizable. They are the two parts of the proposed benchmark: the evaluation of ROI encryption fineness and ROI region perturbation effect.
V-A The Evaluation of ROI Encryption Fineness
@@
V-B The Evaluation of ROI Region Perturbation Effect
@@
V-B1 Analysis of Subjective Vision
To assess the perceptual impact of encryption on the ROI region, we conduct subjective vision comparison by presenting encrypted video frames to the observers. This leverages participant feedback to evaluate the level of perturbation in the ROI region. Specifically, at least two video sequences with different resolutions and texture complexities are selected. For each video sequence, we choose the first and last frames, and show their original frames, encrypted frames (the proposed algorithm), and the corresponding decrypted frames (the proposed algorithm), which can refer to Fig. 5. Subsequently, the whole encrypted frames, encrypted images of ROI region, and encrypted edge images of ROI region are compared with those produced by State-of-the-Art algorithms, which can refer to Fig. 7, Fig. 8, and Fig. 9. Furthermore, various subjective and intuitive visualization examples can be included in the future to illustrate the encryption effectiveness of the proposed algorithm.
V-B2 Analysis of Objective Indicators
Unlike global encryption algorithms, evaluating an ROI encryption algorithm presents unique challenges, as it encrypts only the ROI region of a video frame, and the extracted ROI region may vary across different algorithms. Consequently, the objective indicators used for global encryption cannot be directly applied. To address this issue, we designed an accurate ROI region extraction criterion. This criterion enables precise extracting of all ROI regions based on different ROI region recognition algorithms. Subsequently, we apply various objective evaluation indicators to these regions to assess the security of the encryption algorithm.
ROI Region Extraction Criterion. In video coding standards, video data are typically compressed at the level of coding units rather than individual pixels. Therefore, to accurately assess the perturbation effect within the ROI region, evaluation metrics should be applied to the actual coding units that have been encrypted. To this end, we propose an ROI region extraction criterion based on basic coding units, enabling precise identification of all encrypted units and ensuring that the evaluation results correspond strictly to the actual encrypted regions.
Definition of Basic Coding Unit
We first define $U$ as a “basic coding unit” in video coding. The basic coding unit is a general concept, but its specific form varies across different video coding standards. For instance, it corresponds to a Macroblock (MB) in H.264/AVC and to a Coding Unit (CU) in H.265/HEVC.
Identification of Target Unit Set
For any video frame, let $S_{\text{all}}$ denote the set of all basic coding units in the frame:
@@ V-B2 Analysis of Objective Indicators
Data Extraction
To compute objective evaluation metrics, two types of data are extracted according to $S_{\text{ROI}}$:
@@
Objective Evaluation Indicators. After extracting the ROI region of each frame based on the proposed ROI region extraction criterion, we can integrate a comprehensive set of objective evaluation indicators to systematically assess the perturbation effect of ROI region. These indicators include: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Edge Difference Ratio (EDR), information entropy, Pixel Change Rate (NPCR), Unified Average Change Intensity (UACI), and Bit Rate Change. To provide a clear overview, Table III summarizes the key objective indicators along with their calculation formulas and evaluation dimensions. The following provides a detailed interpretation of the indicators.
TABLE III: Video encryption objective evaluation indicators
@@
Information entropy is used to measure the randomness and uncertainty of encrypted video. The closer the information entropy value is to the ideal value of 8, the stronger the randomness and uncertainty of the encrypted video. In the formula, $L$ represents the binary representation length of the pixel value, and $p(I_{j})$ is the probability of the pixel value $I_{j}$ appearing.
@@
In summary, the objective indicators in the proposed benchmark measure the performance in terms of video quality, security strength, and compression efficiency, helping researchers effectively identify the advantages and disadvantages of various ROI encryption schemes on a standardized platform. In the future, more objective indicators can be added to enrich the proposed benchmark.
@@ VI-A Experimental Setup
The proposed ROI selective encryption scheme is implemented in the H.265/HEVC reference software HM 16.9. The experimental computer configuration is as follows: Intel (R) Core (TM) i7-6700 HQ, 2.6 GHz, 16 GB memory, with Windows 10 operating system, Microsoft Visual Studio 2010, MATLAB 2021a and OpenCV 2.4.7 installed. We use a total of 5 YUV sequences as shown in TABLE IV as our test videos, with resolutions ranging from $352\times 288$ to $2560\times 1600$, encoded at a frame rate of 60 fps, using the profile “encoder_lowdelay_main”, and the Group of Pictures (GOP) is “IBBB”. The QP is set to 8, 24 and 40 respectively. the visual perception network used in our ROI region recognition module is YOLOv8 [wang2024dph] in the experiment. The three-level encryption strategy proposed in this paper are compared with two latest H.265/HEVC ROI encryption schemes, including Yu [yu2023coding] and Taha [taha2018end]. we encrypt 50 frames of each sequence by the three schemes for comparison.
TABLE IV: YUV test sequence
@@
Figure 4: Detection results before and after training
@@ VI-B Examples
Some sample frames of the proposed three-level encryption strategy on the Akiyo video sequence are shown in Fig. 5. Specifically, the frame #1 and #50 of the video sequence are selected to intuitively demonstrate the encryption and decryption effects. The experimental results show that our encryption scheme performs effectively in H.265/HEVC codec. The encrypted frames show severe visual distortion in ROI region, and we can successfully recover the original Akiyo video sequence from the encrypted bitstream.
Figure 5: Some examples of proposed strategy on Akiyo sequence
@@ VI-C Performance Analysis
To evaluate the effectiveness of the proposed scheme, this section presents a comprehensive performance analysis conducted based on the proposed benchmark in Section V. The evaluation is structured into two components: ROI region fineness and ROI region perturbation effect.
@@ VI-C2 Comparison of ROI Region Perturbation Effect
Comparison of Subjective Vision. Fig. 7 shows the frame #1 and #50 of the Johnny sequence encrypted by five encryption strategies under $QP=24$. Meanwhile, In order to present the distortion difference more clearly, we further slice the ROI region and make a detailed comparison of the distortion effect in Fig. 8. It can be clearly seen from Fig. 8 that the visual distortion of Fig. 8(f) and Fig. 8(l) are the most serious, especially the texture details and contours of the face are almost unrecognizable. This is because the advanced level encryption uses edge scrambling in addition to the encryption syntax elements, causing greater distortion. However, since Yu[yu2023coding] and Taha[taha2018end] both encrypt the IPM, their encryption performance is better than that of the basic level encryption but worse than our enhanced level encryption. The reason is that enhanced level encryption introduces the encryption of Delta QP and MVPIdx on the basis of basic level encryption which can produce stronger visual distortion.
@@
Comparison of Objective Indicators. According to TABLE III, the objective indicators analysis includes: PSNR, SSIM, EDR, information entropy, NPCR, UACI and Bit rate change.
@@
The testing results of PSNR and SSIM are recorded in TABLE V and TABLE VI, where the optimal results are marked in bold and the suboptimal results are underlined. The results in the following tables of the other indicators are also highlighted in the same way. As can be seen from the tables, the PSNR and SSIM results of the advanced level encryption are almost the optimal and that of the enhanced level encryption are suboptimal. They are both better than that of Yu [yu2023coding] and Taha [taha2018end].
TABLE V: Average PSNR of the compared strategies under different QP
@@
TABLE VI: Average SSIM of the compared strategies under different QP
@@
we first give the encrypted visual effect in edge region of all the compared strategies in Fig. 9. It can be found that Yu[yu2023coding] and Taha[taha2018end] have large edge distortion than the basic level encryption resulting from the perturbation of IPM. However, both the edge distortion of enhanced and advanced level encryption are better than Yu[yu2023coding] and Taha[taha2018end]. The reason is that the enhanced level encryption encrypts MVPIdx and Delta QP and advanced level encryption scramble the edge coefficients, so the advanced level encryption reaches the best results. Meanwhile, the average values of EDR are also listed in TABLE VII. From the table, the advanced level encryption and the enhanced level encryption achieve the optimal and suboptimal results, which are better than Yu[yu2023coding] and Taha[taha2018end]. The result is consistent with the edge image in Fig 9.
@@
TABLE VII: Average EDR of the compared strategies
@@
the results of the average information entropy of different video sequences in all the compared strategies are shown in Fig. 10. It can be seen that the information entropy of the original video is relatively low, ranging from 6.8 to 7.4, and the information entropy of Yu[yu2023coding] and Taha[taha2018end] is higher. However, the advanced level encryption achieve the highest information entropy, and the result of CLASS_E reached about 7.95, close to the theoretical maximum value of 8.
@@
TABLE VIII: Average NPCR and UACI of the compared strategies
@@
The testing results of NPCR and UACI of all the compared strategies are listed in TABLE VIII. the theoretical values of NPCR and UACI are 99.609% and 33.464%, respectively. Therefore. It can be seen from the table that advanced level encryption obtain the optimal NPCR and UACI, which proves it has good resistance to differential attacks.
@@
The experimental results of bit rate change are shown in TABLE IX. From the analysis in Section III-B, whether the encryption of the syntax element will cause an increase in bit rate depends on its entropy coding mode. Encrypting the syntax elements coded by regular mode will inevitably cause an increase in bit rate, Since Yu[yu2023coding], Taha[taha2018end], the enhanced and advanced level encryption all encrypting the syntax elements coded by regular mode, lead to bit rate increment. However, the basic level encryption only encrypts the syntax elements coded by bypass mode, which getting the best results with a zero bit rate increment.
TABLE IX: Average bit rate change of the compared strategies

### v2

@@ V-C Evaluation Methods
The IoU between the encrypted regions and the detected ROIs can quantify how accurately sensitive areas are covered by the encryption. Meanwhile, the perturbation effect within the extracted ROI regions, which affects whether the content in these regions is sufficiently perturbed to become unrecognizable. They are the two parts of evaluation methods in our proposed benchmark: the evaluation of ROI encryption fineness and ROI region perturbation effect.
V-C1 The Evaluation of ROI Encryption Fineness
@@
V-C2 The Evaluation of ROI Region Perturbation Effect
@@
Analysis of Subjective Vision
To assess the perceptual impact of encryption on the ROI region, we conduct subjective vision comparison by presenting encrypted video frames to the observers. This leverages participant feedback to evaluate the level of perturbation in the ROI region. Specifically, at least two video sequences with different resolutions and texture complexities are selected. For each video sequence, we choose the first and last frames, and show their original frames, encrypted frames (the proposed algorithm), and the corresponding decrypted frames (the proposed algorithm), which can refer to Fig. 5. Subsequently, the whole encrypted frames, encrypted images of ROI region, and encrypted edge images of ROI region are compared with those produced by SOTA comparative algorithms, which can refer to Fig. 7, Fig. 8, and Fig. 9. Furthermore, various subjective and intuitive visualization examples can be included in the future to illustrate the encryption effectiveness of the proposed algorithm.
Analysis of Objective Indicators
Unlike global encryption algorithms, evaluating an ROI encryption algorithm presents unique challenges, as it encrypts only the ROI region of a video frame, and the extracted ROI region may vary across different algorithms. Thus, the objective indicators used for global encryption cannot be directly applied. To address this issue, we designed an accurate ROI region extraction criterion. This criterion enables precise extracting of all ROI regions based on different ROI region recognition algorithms. Then, we apply various objective evaluation indicators to these regions to assess the security of the encryption algorithm.
ROI Region Extraction Criterion. In video coding standards, videos are typically compressed at the level of coding units rather than individual pixels. Therefore, to accurately assess the perturbation effect within the ROI region, evaluation metrics should be applied to the actual coding units that have been encrypted. To this end, we propose an ROI region extraction criterion based on basic coding units, enabling precise identification of all encrypted units and ensuring that the evaluation results correspond strictly to the actual encrypted regions.
Definition of Basic Coding Unit: We first define $U$ as a “basic coding unit” in video coding. The basic coding unit is a general concept, it represents the basic block structure of video encoding. Meanwhile, its specific form varies across different video coding standards. For instance, it corresponds to a Macroblock (MB) in H.264/AVC and to a Coding Unit (CU) in H.265/HEVC.
Identification of Target Unit Set: For any video frame, let $S_{\text{all}}$ denote the set of all basic coding units in the frame:
@@
Data Extraction: To compute objective evaluation metrics, two types of data are extracted according to $S_{\text{ROI}}$:
@@
Objective Evaluation Indicators. After extracting the ROI region of each frame based on the proposed ROI region extraction criterion, we can integrate a comprehensive set of objective evaluation indicators to systematically assess the perturbation effect of ROI region. These indicators include: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), Edge Difference Ratio (EDR), information entropy, Pixel Change Rate (NPCR), Unified Average Change Intensity (UACI), and Bit Rate Change. To provide a clear overview, TABLE V summarizes the key objective indicators along with their calculation formulas and evaluation dimensions. The following provides a detailed interpretation of the indicators.
TABLE V: Video encryption objective evaluation indicators
@@
Information entropy is used to measure the randomness and uncertainty of encrypted video. The closer the information entropy value is to the ideal value of 8, the stronger the randomness and uncertainty of the encrypted video. In the formula, $L$ is the binary representation length of the pixel value, $p(I_{j})$ is the probability of the pixel value $I_{j}$ appearing.
@@
In summary, the seven objective indicators in the proposed benchmark measure the performance in terms of video quality, security strength, and compression efficiency, helping researchers effectively identify the advantages and disadvantages of various ROI encryption schemes on a standardized platform. In the future, more objective indicators can be added to enrich the proposed benchmark.
@@ VI-A Experimental Setup
The proposed ROI selective encryption scheme is implemented in the H.265/HEVC reference software HM 16.9. The experimental computer configuration is as follows: Intel (R) Core (TM) i7-6700 HQ, 2.6 GHz, 16 GB memory, with Windows 10 operating system, Microsoft Visual Studio 2010, MATLAB 2021a and OpenCV 2.4.7 installed. The visual perception network used in our ROI region recognition module is YOLOv8 [wang2024dph] in the experiment.
Figure 4: Detection results before and after training
The video sequences selected for the experiment are taken from the Xiph.org Media in TABLE III of the proposed benchmark. Five sequences containing facial information were chosen for testing. The selected YUV video sequences are listed in TABLE VI, with resolutions ranging from $352\times 288$ to $2560\times 1600$, a frame rate of 60 fps, using the “encoder_lowdelay_main” profile, and a GOP structure of “IBBB”. The QP is set to 8, 24, and 40, respectively. As shown in TABLE IV, our benchmark provides six comparison algorithms. Since our algorithm is a ROI SE encryption algorithm for H.265/HEVC, we choose the two H.265/HEVC ROI encryption algorithms Yu [yu2023coding] and Taha [taha2018end] from TABLE IV for comparison. We encrypt 50 frames of each sequence by the three schemes for comparison.
TABLE VI: YUV test sequence
@@ VI-B Examples
According to V-C2, some sample frames of the proposed three-level encryption strategy in the Akiyo video sequence are shown in Fig. 5.
Figure 5: Some examples of proposed strategy in Akiyo sequence
As described in our proposed benchmark, the frame #1 and #50 of the video sequence are selected to intuitively demonstrate the encryption and decryption effects. The experimental results show that our encryption scheme performs effectively in H.265/HEVC codec. The encrypted frames show severe visual distortion in ROI region, and we can successfully recover the original Akiyo video sequence from the encrypted bitstream.
@@ VI-C Performance Analysis
To evaluate the effectiveness of the proposed scheme, this section presents a comprehensive performance analysis conducted based on the proposed benchmark in Section V. The evaluation is structured into two components: ROI encryption fineness and ROI region perturbation effect.
@@ VI-C2 Comparison of ROI Region Perturbation Effect
Comparison of Subjective Vision. According to V-C2, the frame #1 and #50 of the Johnny sequence encrypted by five encryption strategies under $QP=24$ are listed in Fig. 7. Meanwhile, In order to present the distortion difference more clearly, we further slice the ROI region and make a detailed comparison of the distortion effect in Fig. 8. It can be clearly seen from Fig. 8 that the visual distortion of Fig. 8(f) and Fig. 8(l) are the most serious, especially the texture details and contours of the face are almost unrecognizable. This is because the advanced level encryption uses edge scrambling in addition to the encryption syntax elements, causing greater distortion. However, since Yu[yu2023coding] and Taha[taha2018end] both encrypt the IPM, their encryption performance is better than that of the basic level encryption but worse than our enhanced level encryption. The reason is that enhanced level encryption introduces the encryption of Delta QP and MVPIdx on the basis of basic level encryption which can produce stronger visual distortion. Meanwhile, it can be seen from Fig.8 that the ROI region detected by the Taha[taha2018end] is coarser than ours, containing more background areas, consistent with the results in Fig.6.
@@
Comparison of Objective Indicators. According to TABLE V, we evaluate the objective disturbance effects from the seven indicators: PSNR, SSIM, EDR, information entropy, NPCR, UACI and Bit rate change.
@@
The testing results of PSNR and SSIM are recorded in TABLE VII and TABLE VIII, where the optimal results are marked in bold and the suboptimal results are underlined. The results in the following tables of the other indicators are also highlighted in the same way. As can be seen from the tables, the PSNR and SSIM results of the advanced level encryption are almost the optimal and that of the enhanced level encryption are suboptimal. They are both better than that of Yu [yu2023coding] and Taha [taha2018end].
TABLE VII: Average PSNR of the compared strategies under different QP
@@
TABLE VIII: Average SSIM of the compared strategies under different QP
@@
we first give the encrypted visual effect in edge region of all the compared strategies in Fig. 9. It can be found that Yu[yu2023coding] and Taha[taha2018end] have large edge distortion than the basic level encryption resulting from the perturbation of IPM. However, both the edge distortion of enhanced and advanced level encryption are better than Yu[yu2023coding] and Taha[taha2018end]. The reason is that the enhanced level encryption encrypts MVPIdx and Delta QP and advanced level encryption scramble the edge coefficients, so the advanced level encryption reaches the best results. Meanwhile, the average values of EDR are also listed in TABLE IX. From the table, the advanced level encryption and the enhanced level encryption achieve the optimal and suboptimal results, which are better than Yu[yu2023coding] and Taha[taha2018end]. The result is consistent with the edge image in Fig 9.
@@
TABLE IX: Average EDR of the compared strategies
@@
the results of the average information entropy of different video sequences in all the compared strategies are shown in Fig. 10. It can be seen that the information entropy of the original video is relatively low, ranging from 6.8 to 7.4, and the information entropy of Yu[yu2023coding] and Taha[taha2018end] is higher. However, the advanced level encryption achieve the highest information entropy, and the result of CLASS_E reached approximately 7.95, close to the theoretical maximum value of 8.
@@
TABLE X: Average NPCR and UACI of the compared strategies
@@
The testing results of NPCR and UACI of all the compared strategies are listed in TABLE X. the theoretical values of NPCR and UACI are 99.609% and 33.464%, respectively. Therefore. It can be seen from the table that advanced level encryption obtain the optimal NPCR and UACI, which proves it has good resistance to differential attacks.
@@
The experimental results of bit rate change are shown in TABLE XI. From the analysis in Section III-B, whether the encryption of the syntax element will cause an increase in bit rate depends on its entropy coding mode. Encrypting the syntax elements coded by regular mode will inevitably cause an increase in bit rate, Since Yu[yu2023coding], Taha[taha2018end], the enhanced and advanced level encryption all encrypting the syntax elements coded by regular mode, lead to bit rate increment. However, the basic level encryption only encrypts the syntax elements coded by bypass mode, which getting the best results with a zero bit rate increment.
TABLE XI: Average bit rate change of the compared strategies